Tara, an 18-year-old, just graduated from high school. She now has three choices:

1)Go to work in retail sales and earn $25,000/year until she retires at age 65.

2)Go to cosmetology school for one year during which time she earns nothing and has to pay $11,000 for school. After she finishes cosmetology school she can get a job in the beauty industry and earn $40,000/yearuntil she retires at age 65.

3)Go to college for four years during which time she earns nothing and has to pay $20,000 per year for school. After she finishes her degree she can get a job as a policy analyst and earn $60,000/year until she retires at age 65. Draw Tara’s three different possible age-earnings profiles on one graph. Label the x-and y-axes, label each profile with the educational choice involved, and label the important wages and ages along the axes. Hint: costs for education should be considered negative wages.
Answer #1

On x axis we have year and y axis the amount is given. If we will construct it on a graph we will get a graph as shown above.

Here the interest rate is not given hence assume i= 10%

First case when she starts working from 18 years. She would be earning $ 25,000/ year till 65 years. Refer the attached picture

The present value of cash flow stream = 25,000(P/A,10\%,47)

= 25,000 × 9.8866

= $ 247,165.45

2. Refer the attached picture

The PV of the cash flow

=(P/F,10%,1) [-11,000 + 40,000(P/A,10%,46)]

= $ 349,101.09

3. Refer the attached picture

PV = -20000(P/A,10%,4)+60000(P/F,10%,4)(P/A,10%,43)

= - 63,397.31 + 403,005.16

= $ 339,607.85

Second option has higher present value thus she must go to college for 1 year then start working.

Please contact if having any query will be obliged to you for your generous support. Please help me it mean alot to me thank you.

Amount 30000 25000 20000 15000 10000 5000 0 19 20 21 22 23 24 25 65

Amount 50000 Cash Flow 40000 30000 20000 10000 -10000 -20000 19 20 21 22 23 24 25 Amount -11000 40000 40000 40000 40000 40000 40000 65 40000

Amount 80000 60000 40000 20000 23 24 65 -20000 -40000